OpenStudy (anonymous):

The prom committee has hired a caterer that charges a $600 service fee and $13 per person. Write the linear equation representing this scenario.

OpenStudy (anonymous):

It's not a choice. These are y = 0.42x + 4.41 y = 6x + 4.41 y = 0.42x + 1.89 y = 6x + 1.89

OpenStudy (anonymous):

Yeah your y = was wrong.